How would I solve 16x^2=8x-1.
Answer by stanbon(75887)

solve 16x^2=8x-1
-----------------
16x^2-8x+1 = 0
(4x-1)(4x-1) = 0
---
x = 1/4 with multiplicity 2.
============================
Cheers,
Stan H.
==========